# Optimal Control of an SIR Model with Noncompliance as a Social Contagion

### v1

@@ Abstract.
We propose and study a compartmental model for epidemiology with human behavioral effects. Specifically, our model incorporates governmental prevention measures aimed at lowering the disease infection rate, but we split the population into those who comply with the measures and those who do not comply and therefore do not receive the reduction in infectivity. We then allow the attitude of noncompliance to spread as a social contagion parallel to the disease. We derive the reproductive ratio for our model and provide stability analysis for the disease-free equilibria. We then propose a control scenario wherein a policy-maker with access to control variables representing disease prevention mandates, treatment efforts, and educational campaigns aimed at encouraging compliance minimizes a cost functional incorporating several cost concerns. We characterize optimal controls via the Pontryagin optimality principle and present simulations which demonstrate the behavior of the control maps in several different parameter regimes.
@@ 1. Introduction
Incorporation and assessment of the effects of heterogeneous behavior into epidemiological models is becoming increasingly popular (see, for example, [4, 21, 27, 34, 36], or [43] where the authors propose a rather general framework similar to the modeling we use below, though with linear transfer between behavior classes). Likewise, there are several recent studies of optimal control for epidemiological models in different modeling scenarios, such as optimal treatment or vaccination [6, 13, 26, 50], optimal quarantine under a variety of limitations [3, 15], and assessment of the effects of contact heterogeneity [14] or the use of different types of cost functionals [44]. However, very few studies include both optimal control for epidemiology and mechanistic modeling of human behavior. One exception is [8] which takes a more theoretical look at a system of controlled partial differential equations describing an epidemic. Other studies incorporate control and human behavior more implicitly, by, for example, allowing infection rates to depend on ambient dynamics under the assumption that people will change their behavior based on local disease effects [37] or media coverage [18].
In this manuscript, we develop a compartmental epidemiological model which incorporates human behavior in an explicit, mechanistic manner. We then consider optimal control of the model and observe the differences in optimal control strategies in different parameter regimes. The remainder of the manuscript is organized as follows. In section 2, we introduce our model, derive the reproductive ratio, and discuss stability of disease-free equilibria (though a full proof of our stability theorem is relegated to appendix A). In section 3, we introduce the optimal control problem that we will solve and give a characterization of the optimal control maps via the Pontryagin optimality principle. In section 4, we discuss the results of simulation of our model in different parameter regimes. Finally, section 5 includes a brief conclusion and discussion of avenues for future work.
@@ 2. An SIR Model with Behavioral Effects: Disease-Free Equilibria and Reproductive Ratios
(2.2) $\begin{split}\frac{dS}{dt}&=(1-\xi)b-\beta(1-\alpha(t))S(I+I^{*})-(\overline{\mu}-\mu(t))SN^{*}+\nu(t)S^{*}-\delta S,\\ \frac{dI}{dt}&=\beta(1-\alpha(t))S(I+I^{*})-(\gamma+\eta(t))I-(\overline{\mu}-\mu(t))IN^{*}+\nu(t)I^{*}-\delta I,\\ \frac{dR}{dt}&=(\gamma+\eta(t))I-(\overline{\mu}-\mu(t))RN^{*}+\nu(t)R^{*}-\delta R,\\ \frac{dS^{*}}{dt}&=\xi b-\beta S^{*}(I+I^{*})+(\overline{\mu}-\mu(t))SN^{*}-\nu(t)S^{*}-\delta S^{*},\\ \frac{dI^{*}}{dt}&=\beta S^{*}(I+I^{*})-\gamma I^{*}+(\overline{\mu}-\mu(t))IN^{*}-\nu(t)I^{*}-\delta I^{*},\\ \frac{dR^{*}}{dt}&=\gamma I^{*}+(\overline{\mu}-\mu(t))RN^{*}-\nu(t)R^{*}-\delta R^{*}.\end{split}$
Here the birth rate $b$, death rate $\delta$, infectivity rate $\beta$ and recovery rate $\gamma$ have the same interpretation as in (2.1). New parameters are as follows: $\xi\in[0,1]$ is the probability with which newly introduced members of the population are noncompliant, and hence the birth rate $b$ is split between the $S$ and $S^{*}$ equations with contributions $(1-\xi)b$ and $\xi b$ respectively. The total noncompliant population is defined by $N^{*}=S^{*}+I^{*}+R^{*}$ and is involved in new mass action terms which cause transfer from $S,I,R$ to $S^{*},I^{*},R^{*}$ with baseline infectivity rate $\overline{\mu}$ (the assumption being that anytime a compliant individual comes in contact with a noncompliant individual, they have some chance of becoming noncompliant). The remaining parameters $\alpha,\eta,\mu,\nu$ represent the policy-maker’s attempts to stop the spread of the disease and/or noncompliance. While we will typically suppress their $t$-dependence, we have written them as functions of $t$ in (2.2) to emphasize that these will be treated as control variables that the policy-maker can adjust throughout the course of the epidemic. Their respective interpretations and admissible values are:
@@
    $\alpha(\cdot)\in[0,1]$ - reduction in infectivity among interactions with the compliant susceptible population.
@@
    $\mu(\cdot)\in[0,\overline{\mu}]$ - reduction in spread of noncompliance due to public health information initiatives.
@@
The uncontrolled scenario would be $\alpha(\cdot),\eta(\cdot),\mu(\cdot),\nu(\cdot)\equiv 0$, and the maximally controlled scenario would have each control variable identically equal to its maximum value. Note that we do not consider vaccination, which can be modeled as either an ordinary or “singular” control variable as discussed in [26], opting instead to focus on the early stages of infection when vaccination is unavailable and NPIs play a larger role. In this way, our model is similar to the ones analyzed in [7] in the ODE setting, [39] in the stochastic ODE setting, and [8, 38] in the PDE setting. However, [7, 38, 39] do not consider the optimal control problem, and the work in [8] does not incorporate treatment as a control, and is also more focused on analytic results regarding existence and characterization of solutions to a PDE-constrained optimization problem. Here we are interested in empirically observing features of the optimal control maps in different parameter regimes.
In the ensuing section, we will introduce a cost functional for the policy-maker, and evaluate optimal control strategies. Before that, we cover some standard considerations regarding compartmental epidemiological models; namely, nonnegativity of solutions, derivation of the reproductive ratio, and stability of disease-free equilibria. For the remainder of this section, we treat $\alpha,\eta,\mu,\nu$ as fixed, nonnegative, constant values. We will go through this quickly; much of the same work is included in more detail in [39], albeit for a slightly different system of equations.
Because we will use the next generation matrix method [19, 46, 47] to derive the reproductive ratio for our system, we write (2.1) abstractly as
(2.3) $x^{\prime}=\mathcal{F}(x)+\mathcal{V}^{+}(x)-\mathcal{V}^{-}(x)$
where we reorder the variables $x=(I,I^{*},S,R,S^{*},R^{*})$ by bringing the infected compartments to the front, and define
(2.4) $\mathcal{F}(x)=\begin{bmatrix}\beta(1-\alpha)S(I+I^{*})\\ \beta S^{*}(I+I^{*})\\ 0\\ 0\\ 0\\ 0\end{bmatrix}$
and
(2.5) $\mathcal{V}^{+}(x)=\begin{bmatrix}\nu I^{*}\\ (\overline{\mu}-\mu)IN^{*}\\ (1-\xi)b+\nu S^{*}\\ (\gamma+\eta)I+\nu R^{*}\\ \xi b+(\overline{\mu}-\mu)SN^{*}\\ \gamma I^{*}+(\overline{\mu}-\mu)RN^{*}\end{bmatrix},\,\,\,\,\,\,\,\,\,\,\mathcal{V}^{-}(x)=\begin{bmatrix}(\gamma+\eta+\delta+(\overline{\mu}-\mu)N^{*})I\\ (\gamma+\nu+\delta)I^{*}\\ (\beta(1-\alpha)(I+I^{*})+(\overline{\mu}-\mu)N^{*}+\delta)S\\ ((\overline{\mu}-\mu)N^{*}+\delta)R\\ (\beta(I+I^{*})+\nu+\delta)S^{*}\\ (\nu+\delta)R^{*}\end{bmatrix}$.
Here $\mathcal{F}$ includes all terms which correspond to adding new infections into the system, $\mathcal{V}^{+}$ includes all terms corresponding to in-flow to a compartment caused by anything other than new infections, and $\mathcal{V}^{-}$ includes all terms corresponding to outflow from a compartment for any reason. We note in particular that $\mathcal{F},\mathcal{V}^{+},\mathcal{V}^{-}\geq 0$ when $S,I,R,S^{*},I^{*},R^{*}\geq 0$.
@@
Assuming nonnegative initial conditions, (2.2) admits a unique, global-in-time, nonnegative solution.
Proof.
A unique local solution to (2.2) exists as consequence of the Picard-Lindelöf theorem since the nonlinearities are Lipschitz. Nonnegativity is preserved as long as the solution exists because, looking at each component $x_{j}$ of (2.3), we have $x_{j}^{\prime}\geq 0$ when $x_{j}=0$ since the $i^{\text{th}}$ component of $\mathcal{V}^{-}$ is proportional to $x_{i}$. Given nonnegativity, we see that solutions cannot blow up because the total population $N_{\text{total}}=S+I+R+S^{*}+I^{*}+R^{*}$ satisfies
@@
which serves as an upper bound on each of the subpopulations. ∎
@@
We can assume without loss of generality that $N_{\text{total}}(0)=1$, whereupon alternately setting $\tfrac{b}{\delta}\geq 1$ or $\tfrac{b}{\delta}<1$ would correspond to situations where the population is increasing or where the population is strictly decreasing, respectively. For ease of analysis, we will always assume that $\tfrac{b}{\delta}\geq 1=N(0)$, so that $\tfrac{b}{\delta}$ serves as a bound on the total population.
A disease-free equilibrium (DFE) point for (2.2) is an equilibrium point wherein $I=I^{*}=0$. By adding the equations for $R,R^{*}$ together, we see that any DFE will also have $R=R^{*}=0$, whereupon $x=(0,0,s,0,s^{*},0)$ is a DFE if and only if
(2.7) $\begin{split}(1-\xi)b-(\overline{\mu}-\mu)ss^{*}+\nu s^{*}-\delta s&=0,\\ \xi b+(\overline{\mu}-\mu)ss^{*}-\nu s^{*}-\delta s^{*}&=0.\end{split}$
Because we are looking for physically meaningful solutions, we only consider values $0\leq s,s^{*}\leq\frac{b}{\delta}$. With this restriction in mind, we specify the different parameter regimes which lead to existence of different DFEs. If we assume $\xi=0$ so that all newly introduced members of the population are compliant, there are two solutions which can be physically meaningful:
(2.8) $\begin{split}s_{1}&=\tfrac{b}{\delta},\\ s^{*}_{1}&=0,\end{split}\hskip 56.9055pt\text{ and }\hskip 56.9055pt\begin{split}s_{2}&=\tfrac{\nu+\delta}{\overline{\mu}-\mu},\\ s^{*}_{2}&=\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu},\end{split}$
@@
The first is always physically meaningful and represents the “best case scenario” for the policy-maker wherein the whole population is compliant. The second is physically meaningful and distinct only if $\tfrac{b}{\delta}>\frac{\nu+\delta}{\overline{\mu}-\mu}.$ Since $\mu\in[0,\overline{\mu}]$, the denominator of $\frac{\nu+\delta}{\overline{\mu}-\mu}$ is nonnegative, so in the case that $\mu\to\overline{\mu}$, the second DFE is no longer meaningful.
@@
If instead $\xi\in(0,1]$ so that some portion of the newly introduced population is noncompliant, (2.7) has a unique physically meaningful solution
(2.9) $\begin{split}s_{3}&=\tfrac{1}{2}\left(\tfrac{b}{\delta}+\tfrac{\delta+\nu}{\overline{\mu}-\mu}-\sqrt{\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\overline{\mu}-\mu}\right)^{2}+\tfrac{4\xi b}{\overline{\mu}-\mu}}\right),\\ s^{*}_{3}&=\tfrac{1}{2}\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\overline{\mu}-\mu}+\sqrt{\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\overline{\mu}-\mu}\right)^{2}+\tfrac{4\xi b}{\overline{\mu}-\mu}}\right).\end{split}$
@@
In the limit as $\xi\to 0$, we find $(s_{3},s_{3}^{*})\to(s_{1},s_{1}^{*})$ when $\tfrac{b}{\delta}\leq\frac{\nu+\delta}{\overline{\mu}-\mu}$, and $(s_{3},s_{3}^{*})\to(s_{2},s_{2}^{*})$ when $\tfrac{b}{\delta}>\frac{\nu+\delta}{\overline{\mu}-\mu}$. A special case of (2.9) is the “worst case scenario” for the policy-maker wherein $\xi=1$ and $\nu=0$, meaning that all newly introduced members of the population are noncompliant and noncompliance is a permanent state. In this case, the DFE is $s_{3}=0,s^{*}_{3}=\tfrac{b}{\delta}$.
Given this, we can compute the reproductive ratio $\mathscr{R}_{0}(s,s^{*})$ corresponding to the DFE $x_{s,s^{*}}=(s,s^{*})$. Roughly speaking, $\mathscr{R}_{0}(s,s^{*})$ can be interpreted as the expected number of secondary infections that result from any one infection. This can be computed via the next generation matrix method [19, 46, 47]: it is the spectral radius of the matrix $FV^{-1}$ where $F$ and $V$ are, respectively, the leading principal submatrices of $D\mathcal{F}(x_{s,s^{*}})$ and $D\mathcal{V}^{-}(x_{s,s^{*}})-D\mathcal{V}^{+}(x_{s,s^{*}})$ corresponding to the infected compartments, where $\mathcal{F},\mathcal{V}^{+},\mathcal{V}^{-}$ are as in (2.4),(2.5). In our case, these are
(2.10) $F=\begin{bmatrix}\beta(1-\alpha)s&\beta(1-\alpha)s\\ \beta s^{*}&\beta s^{*}\end{bmatrix},\hskip 56.9055ptV=\begin{bmatrix}\gamma+\eta+\delta+(\overline{\mu}-\mu)s^{*}&-\nu\\ -(\overline{\mu}-\mu)s^{*}&\gamma+\nu+\delta\end{bmatrix}$.
Because $F$ is rank 1, the spectral radius $\rho(FV^{-1})=\mathscr{R}_{0}(s,s^{*})$ can easily be computed by hand. We write it in two different equivalent ways to point out behavior with respect to changes in different parameters:
(2.11) $\begin{split}\mathscr{R}_{0}(s,s^{*})&=\frac{\beta\Big{[}(1-\alpha)\Big{(}\gamma+\nu+\delta+(\overline{\mu}-\mu)s^{*}\Big{)}s+\Big{(}\gamma+\eta+\nu+\delta+(\overline{\mu}-\mu)s^{*}\Big{)}s^{*}\Big{]}}{(\gamma+\delta)(\gamma+\eta+\nu+\delta+(\overline{\mu}-\mu)s^{*})+\eta\nu}\\ &=\frac{\beta\Big{[}(1-\alpha)\Big{(}1-\tfrac{\eta}{\gamma+\eta+\nu+\delta+(\overline{\mu}-\mu)s^{*}}\Big{)}s+s^{*}\Big{]}}{\gamma+\delta+\frac{\eta\nu}{\gamma+\eta+\nu+\delta+(\overline{\mu}-\mu)s^{*}}}.\end{split}$
@@
The second representation lends itself to a reasonably natural interpretation, with the main complicating factor being the difference in recovery rates between the compliant and noncompliant populations due to the compliant population seeking treatment represented by $\eta$. In the worst case scenario, where $\xi=1,\nu=0$ so that the DFE is $(s,s^{*})=(0,\tfrac{b}{\delta})$, we find $\mathscr{R}_{0}(0,\tfrac{b}{\delta})=\frac{b}{\delta}\cdot\frac{\beta}{\gamma+\delta}.$ This is the reproductive ratio for the basic SIR model (2.1), which signifies that in this worst case scenario, the policy-maker’s control efforts do nothing to stop the spread of the disease. In the best case scenario, $\xi=0$ and $\tfrac{b}{\delta}\leq\frac{\nu+\delta}{\overline{\mu}-\mu}$ so that the only DFE is $(s,s^{*})=(\tfrac{b}{\delta},0)$ and we find that $\mathscr{R}_{0}(\tfrac{b}{\delta},0)=\tfrac{b}{\delta}\cdot\frac{\beta(1-\alpha)}{\gamma+\eta+\delta}.$ This decreases the reproductive ratio when compared with the basic SIR model in two ways: $(1-\alpha)$ in the numerator results from the reduction in infectivity among the compliant susceptible population and the $\eta$ in the denominator results from the increased recovery rate among the compliant infected population. In general, $\mathscr{R}_{0}(s,s^{*})$ can then be viewed as a very complicated interpolation between these two endpoints which is illuminated a bit in the second line of (2.11) where the terms in front of $s$ in the numerator can be seen to represent reduction in disease spread among the compliant population, and that $s^{*}$ has no such modifications represents inefficacy of control efforts among the noncompliant population.

### v2

@@ Abstract.
We propose and study a compartmental model for epidemiology with human behavioral effects. Specifically, our model incorporates governmental prevention measures aimed at lowering the disease infection rate, but we split the population into those who comply with the measures and those who do not comply and therefore do not receive the reduction in infectivity. We then allow the attitude of noncompliance to spread as a social contagion parallel to the disease. We derive the reproductive ratio for our model and provide stability analysis for the disease-free equilibria. We then propose an optimal control scenario wherein a policy-maker with access to control variables representing disease prevention mandates, treatment efforts, and educational campaigns aimed at encouraging compliance minimizes a cost functional incorporating several cost concerns. Via careful analysis of the control-to-state map, we are able to prove existence of optimal controls. Our proof applies to dynamics which can be nonlinear in the control variables and general cost functionals including the case of $L^{1}$ control costs. We numerically resolve optimal strategies using the sequential quadratic Hamiltonian method, a relatively new numerical method for optimal control which is easy to implement and has good convergence theory, as we demonstrate. We test our model in several parameter regimes with specific interest in observing how the policy-maker’s optimal strategies depend on their particular preferences which are expressed via design of different cost functionals.
@@ 1. Introduction
Incorporation and assessment of the effects of heterogeneous behavior into epidemiological models is becoming increasingly popular (see, for example, [5, 26, 32, 39, 41], or [50] where the authors propose a rather general framework similar to the modeling we use below, though with linear transfer between behavior classes). Likewise, there are several recent studies of optimal control for epidemiological models in different modeling scenarios, such as optimal treatment or vaccination [1, 7, 23, 31, 59], optimal quarantine under a variety of limitations [4, 17], and assessment of the effects of contact heterogeneity [16] or the use of different types of cost functionals [51]. However, very few studies include both optimal control for epidemiology and mechanistic modeling of human behavior. One exception is [9] which takes a more theoretical look at a system of controlled partial differential equations describing an epidemic. Other studies incorporate control and human behavior more implicitly, by, for example, allowing infection rates to depend on ambient dynamics under the assumption that people will change their behavior based on local disease effects [42] or media coverage [20].
1.1. Our Contribution.
We develop a compartmental epidemiological model which incorporates human behavior in an explicit, mechanistic manner. We then consider optimal control of the model and observe the differences in optimal control strategies in different parameter regimes. In particular, we prove existence of optimal controls in a generic scenario wherein (i) controls appear in the model nonlinearly, (ii) control costs are not required to be smooth or strongly convex (e.g. $L^{1}$ control costs are allowable), and (iii) the admissible control set contains discontinuous control maps. With small tweaks, our existence proof will go through for many biological models where (i)-(iii) are physically preferable, but are commonly relaxed (e.g. by the restriction to control affine systems and use of $L^{2}$ control costs) for ease of analysis. We resolve optimal controls via the sequential quadratic Hamiltonian (SQH) method [10]. This is relatively new algorithm for optimal control which is black-box free and has a reasonable convergence theory, as demonstrated below. This distinguishes SQH from open source large scale optimization packages (used, for example, in [7, 17]) which are explainable to experts, but often operate as black-boxes for the end user, and also from the popular forward-backward sweep method (used, for example, in [1, 20, 23]) which is easy to implement but does not have a robust convergence theory. To the authors’ knowledge, this method has not yet appeared in the epidemiology literature, though it does appear in the preprint [43] in the context of control for a partial differential equation model for epidemics. Thus, the contribution of this manuscript goes further than the particular model we analyze, but also includes a general existence framework for optimal controls that applies to biologically realistic models, and the introduction to the epidemiology literature of a numerical method for optimal control which may be preferable to those commonly used. The remainder of the manuscript is organized as follows. In section 2, we introduce our model, derive the reproductive ratio, and discuss stability of disease-free equilibria. In section 3, we introduce the optimal control problem that we will solve and prove existence of optimal controls for a reasonably generic cost functional. In particular, our existence proof allows for $L^{1}$ control cost and discontinuous controls, which are physically meaningful for our model. In section 4, we describe the application of the SQH method [10] to our problem, and prove convergence of the algorithm. In section 5, we discuss the results of simulations of our model in different parameter regimes. Specifically, we make some qualitative observations regarding the optimal control strategies that result from different policy-maker preferences. Finally, section 6 includes a brief conclusion and discussion of avenues for future work.
@@ 2. An SIR Model with Behavioral Effects: Disease-Free Equilibria and Reproductive Ratios
(2.2) $\begin{split}\frac{dS}{dt}&=(1-\xi)b-\beta(1-\alpha(t))S((1-\alpha(t))I+I^{*})-(\mu^{*}-\mu(t))SN^{*}+\nu(t)S^{*}-\delta S,\\ \frac{dI}{dt}&=\beta(1-\alpha(t))S((1-\alpha(t))I+I^{*})-(\gamma+\eta(t))I-(\mu^{*}-\mu(t))IN^{*}+\nu(t)I^{*}-\delta I,\\ \frac{dR}{dt}&=(\gamma+\eta(t))I-(\mu^{*}-\mu(t))RN^{*}+\nu(t)R^{*}-\delta R,\\ \frac{dS^{*}}{dt}&=\xi b-\beta S^{*}((1-\alpha(t))I+I^{*})+(\mu^{*}-\mu(t))SN^{*}-\nu(t)S^{*}-\delta S^{*},\\ \frac{dI^{*}}{dt}&=\beta S^{*}((1-\alpha(t))I+I^{*})-\gamma I^{*}+(\mu^{*}-\mu(t))IN^{*}-\nu(t)I^{*}-\delta I^{*},\\ \frac{dR^{*}}{dt}&=\gamma I^{*}+(\mu^{*}-\mu(t))RN^{*}-\nu(t)R^{*}-\delta R^{*}.\end{split}$
Here the birth rate $b$, death rate $\delta$, infectivity rate $\beta$ and recovery rate $\gamma$ have the same interpretation as in (2.1). New parameters are as follows: $\xi\in[0,1]$ is the probability with which newly introduced members of the population are noncompliant, and hence the birth rate $b$ is split between the $S$ and $S^{*}$ equations with contributions $(1-\xi)b$ and $\xi b$ respectively. The total noncompliant population is defined by $N^{*}=S^{*}+I^{*}+R^{*}$ and is involved in new mass action terms which cause transfer from $S,I,R$ to $S^{*},I^{*},R^{*}$ with baseline infectivity rate $\mu^{*}$ (the assumption being that anytime a compliant individual comes in contact with a noncompliant individual, they have some chance of becoming noncompliant). The remaining parameters $\alpha,\eta,\mu,\nu$ represent the policy-maker’s attempts to stop the spread of the disease and/or noncompliance. While we will typically suppress their $t$-dependence, we have written them as functions of $t$ in (2.2) to emphasize that these will be treated as control variables that the policy-maker can adjust throughout the course of the epidemic. Their respective interpretations and admissible values are:
@@
    $\alpha(\cdot)\in[0,\overline{\alpha}]$ - reduction in infectivity or mixing for the compliant population. Here $\overline{\alpha}\in[0,1]$ is the maximum decrease in mixing / infectivity that can be achieved through the use of NPIs. Note, $\alpha(t)$ appears in the mass action terms representing spread of infections; in these terms, the compliant subpopulations $S,I$ are multiplied by $(1-\alpha(t))$, while the noncompliant subpopulations $S^{*},I^{*}$ are not.
@@
    $\mu(\cdot)\in[0,\overline{\mu}]$ - reduction in spread of noncompliance due to public health information initiatives. Here $\overline{\mu}\in[0,\mu^{*}]$ is the maximum achievable decrease in the spread rate for noncompliance.
@@
We define $U_{\text{ad}}=[0,\overline{\alpha}]\times[0,\overline{\eta}]\times[0,\mu^{*}]\times[0,\overline{\nu}]$ to be our set of admissible control actions. The uncontrolled scenario would be $\alpha(\cdot),\eta(\cdot),\mu(\cdot),\nu(\cdot)\equiv 0$, and the maximally controlled scenario would have each control variable identically equal to its maximum value. Note that we do not consider vaccination, which can be modeled as either an ordinary or “singular” control variable as discussed in [31], opting instead to focus on the early stages of infection when vaccination is unavailable and NPIs play a larger role. In this way, our model is similar to the ones analyzed in [8] in the ODE setting, [45] in the stochastic ODE setting, and [9, 44] in the PDE setting. However, [8, 44, 45] do not consider the optimal control problem, and the work in [9] does not incorporate treatment as a control, and is also more focused on analytic results regarding existence and characterization of solutions to a PDE-constrained optimization problem. Here we are interested in empirically observing features of the optimal control maps in different parameter regimes.
In the ensuing section, we will introduce a cost functional for the policy-maker, and evaluate optimal control strategies. Before that, we cover some standard considerations regarding compartmental epidemiological models; namely, nonnegativity of solutions, derivation of the reproductive ratio, and stability of disease-free equilibria. We note that this work is already contained in [45], except in that manuscript $\eta=0$. Because $\eta>0$ causes an asymmetry in the way $I$ and $I^{*}$ are treated, the derivations here are a bit messier, though no more difficult. Thus to avoid duplication, we will cover them tersely and refer the reader to [45] for proofs and expanded reasoning.
@@
Assuming nonnegative initial conditions and $(\alpha,\eta,\mu,\nu)\in L^{\infty}([0,T];U_{\text{ad}})$, (2.2) admits a unique, nonnegative, a.e. solution on the time interval $[0,T].$
The existence and uniqueness follow directly from Caratheodory’s theorem [49, Ch. 1.6]. The preservation of nonnegativity in proposition 2.1 follows since $\frac{dX}{dt}|_{X=0}\geq 0$ whenever all components are nonnegative, for each $X\in\{S,I,R,S^{*},I^{*},R^{*}\}.$ We also note that there is a total population bound. Indeed, defining $N_{\text{total}}=S+I+R+S^{*}+I^{*}+R^{*}$, we see
@@
We can assume without loss of generality that $N_{\text{total}}(0)=1$, whereupon alternately setting $\tfrac{b}{\delta}\geq 1$ or $\tfrac{b}{\delta}<1$ would correspond to situations where the population is increasing or where the population is strictly decreasing.
For the remainder of this section, we treat $\alpha,\eta,\mu,\nu$ as fixed, nonnegative, constant values for the sake of finding disease free equilibrium (DFE) points. The DFEs for out system are equilibrium points $x^{*}=(s,i,r,s^{*},i^{*},r^{*})$ with $i=i^{*}=r=r^{*}=0$ and $s,s^{*}\geq 0$. There are two DFEs if $\xi=0$ (meaning that all newly introduced members of the population are compliant). These are given by
(2.4) $\begin{split}s_{1}&=\tfrac{b}{\delta},\\ s^{*}_{1}&=0,\end{split}\hskip 56.9055pt\text{ and }\hskip 56.9055pt\begin{split}s_{2}&=\tfrac{\nu+\delta}{\mu^{*}-\mu},\\ s^{*}_{2}&=\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\mu^{*}-\mu},\end{split}$
@@
The first is always physically meaningful and represents the “best case scenario” for the policy-maker wherein the whole population is compliant. The second is physically meaningful and distinct only if $\tfrac{b}{\delta}>\frac{\nu+\delta}{\mu^{*}-\mu}.$ Since $\mu\in[0,\mu^{*}]$, the denominator of $\frac{\nu+\delta}{\mu^{*}-\mu}$ is nonnegative, so in the case that $\mu\to\mu^{*}$, the second DFE is no longer meaningful.
@@
If instead $\xi\in(0,1]$ so that some portion of the newly introduced population is noncompliant, there is a unique DFE
(2.5) $\begin{split}s_{3}&=\tfrac{1}{2}\left(\tfrac{b}{\delta}+\tfrac{\delta+\nu}{\mu^{*}-\mu}-\sqrt{\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\mu^{*}-\mu}\right)^{2}+\tfrac{4\xi b}{\mu^{*}-\mu}}\right),\\ s^{*}_{3}&=\tfrac{1}{2}\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\mu^{*}-\mu}+\sqrt{\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\mu^{*}-\mu}\right)^{2}+\tfrac{4\xi b}{\mu^{*}-\mu}}\right).\end{split}$
@@
In the limit as $\xi\to 0$, we find $(s_{3},s_{3}^{*})\to(s_{1},s_{1}^{*})$ when $\tfrac{b}{\delta}\leq\frac{\nu+\delta}{\mu^{*}-\mu}$, and $(s_{3},s_{3}^{*})\to(s_{2},s_{2}^{*})$ when $\tfrac{b}{\delta}>\frac{\nu+\delta}{\mu^{*}-\mu}$. A special case of (2.5) is the “worst case scenario” for the policy-maker wherein $\xi=1$ and $\nu=0$, meaning that all newly introduced members of the population are noncompliant and noncompliance is a permanent state. In this case, the DFE is $s_{3}=0,s^{*}_{3}=\tfrac{b}{\delta}$.
The reproductive ratio $\mathscr{R}_{0}(s,s^{*})$ corresponding to a DFE $x_{s,s^{*}}=(s,0,0,s^{*},0,0)$ can be computed by the next generation matrix method [21, 54, 55]:
(2.6) $\mathscr{R}_{0}(s,s^{*})=\frac{\beta\Big[(1-\alpha)^{2}\Big(1+\tfrac{\alpha\mu s^{*}/(1-\alpha)-\eta}{\gamma+\eta+\nu+\delta+(\mu^{*}-\mu)s^{*}}\Big)s+\Big(1-\tfrac{\alpha\nu}{\gamma+\eta+\nu+\delta+(\mu^{*}-\mu)s^{*}}\Big)s^{*}\Big]}{\gamma+\delta+\frac{\eta\nu}{\gamma+\eta+\nu+\delta+(\mu^{*}-\mu)s^{*}}}.$
@@
This has a reasonably natural interpretation, with the main complicating factor being the difference in recovery rates between the compliant and noncompliant populations due to the compliant population seeking treatment represented by $\eta$. In the worst case scenario, where $\xi=1,\nu=0$ so that the DFE is $(s,s^{*})=(0,\tfrac{b}{\delta})$, we find $\mathscr{R}_{0}(0,\tfrac{b}{\delta})=\frac{b}{\delta}\cdot\frac{\beta}{\gamma+\delta}.$ This is the reproductive ratio for the basic SIR model (2.1), which signifies that in this worst case scenario, the policy-maker’s control efforts do nothing to stop the spread of the disease. In the best case scenario, $\xi=0$ and $\tfrac{b}{\delta}\leq\frac{\nu+\delta}{\mu^{*}-\mu}$ so that the only DFE is $(s,s^{*})=(\tfrac{b}{\delta},0)$ and we find that $\mathscr{R}_{0}(\tfrac{b}{\delta},0)=\tfrac{b}{\delta}\cdot\frac{\beta(1-\alpha)^{2}}{\gamma+\eta+\delta}.$ This decreases the reproductive ratio when compared with the basic SIR model in two ways: $(1-\alpha)^{2}$ in the numerator results from the reduction in infectivity among the compliant populations and the $\eta$ in the denominator results from the increased recovery rate among the compliant infected population. In general, $\mathscr{R}_{0}(s,s^{*})$ can then be viewed as a very complicated interpolation between the two points wherein the compliant susceptible population $s$ is receiving some reduction in infectivity signified by the $(1-\alpha)^{2}$, whereas the noncompliant susceptible population $s^{*}$ receives no such benefit.